A coffee merchant has coffee beans that sell for $9.00 per pound and $12.00 per pound. The two types are to be mixed to create 100 lb of a mixture that will sell for $11.25 per pound. How much of each type of bean should be used in the mixture?
______lbs at $9 per pound
______lbs at $12 per pound

Answer by stanbon(48546)

A coffee merchant has coffee beans that sell for $9.00 per pound and $12.00 per pound. The two types are to be mixed to create 100 lb of a mixture that will sell for $11.25 per pound. How much of each type of bean should be used in the mixture?
----------
Let "x" be the amt of $9 coffee: Value of this is 9x dollars
Then "100-x" is the amt of $12 coffee: Value of this is 12(100-x)=1200-12x dollars
---------
Mixture Data:
Amt of mixture is 100 lbs; Value of this is 1125 dollars.
--------
EQUATION:
value + value = value
9x + 1200-12x = 1125
-3x = - 75
x=25 lbs ( amt of $9 coffee)
100-25=75 lbs ( amt of $12 coffee)
===========
Cheers,
Stan H.